# TURA: Tool-Augmented Unified Retrieval Agent for AI Search

### v1

@@ 4.1.2. Server-level Semantic Index Augmentation
A significant challenge in tool use is the ”lexical gap” between user vernacular and formal API or server descriptions. To bridge this, we perform an extensive offline index augmentation process. For each server $M_{i}\in\mathcal{M}$, we first define its holistic description $D_{i}$. Then, we utilize a generative LLM, $g_{\text{LLM-gen}}$ with a specific prompt (see Appendix A.1), to produce a large, diverse set of synthetic queries , $\mathcal{Q}^{\text{syn}}_{i}$, that a user might plausibly issue to access the functionalities of $M_{i}$.
(3) $\mathcal{Q}^{\text{syn}}_{i}=g_{\text{LLM-gen}}(D_{i},\text{temperature}=T_{\text{high}})$
@@ 4.2. DAG-based Task Planner
The planner, implemented with a highly capable LLM $p_{\text{LLM-plan}}$, is prompted to act as a solution architect with a specific prompt (see Appendix A.3). It analyzes the relationships between sub-queries and the capabilities of the retrieved servers to construct a DAG, $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
@@ 4.3.1. Trajectory Synthesis and Data Curation
This raw data is then subjected to a rigorous, automated curation pipeline. The first stage, Correctness Filtering, employs a judge model, $J_{\text{correct}}$(Huang et al., 2024), to validate each step of a trajectory. This judge scrutinizes for adherence to API schemas, validity of parameter values and logical soundness of the thought process leading to the action. Any trajectory failing these checks is discarded.
Subsequently, the second stage, Efficiency Filtering, uses another judge, $J_{\text{efficient}}$, to analyze the now-correct trajectories for performance. It identifies and flags issues such as action redundancy and path sub-optimality These inefficient trajectories are then either pruned or programmatically corrected. This two-stage curation transforms the noisy expert data into a high-quality, optimal distillation dataset, $D_{\text{distill}}$:
@@
4.3.2. Mixed-Rationale Supervised Fine-Tuning (SFT)
To achieve minimal inference latency, we fine-tune Qwen3(Yang et al., 2025) series, a much smaller model than large-scale LLM like Deepseek-V3 on $\mathcal{D}_{\text{distill}}$ using a mixed-rationale SFT strategy. The training process explicitly leverages the chain-of-thought data. The agent $\mathcal{A}_{\theta}$ is trained to predict the full sequence of tokens, including both the thought and the action. The loss function is the standard cross-entropy loss over the target sequence:
(8) $\mathcal{L}_{\text{SFT}}(\theta)=-\sum_{(st^{\prime},\pi^{\prime})\in\mathcal{D}_{\text{distill}}}\sum_{t=1}^{|\pi^{\prime}|}\log P_{\theta}(y_{t}|y_{<t},st^{\prime})$
@@ 4.3.2. Mixed-Rationale Supervised Fine-Tuning (SFT)
Critically, during online inference, we provide the agent $\mathcal{A}_{\theta}$ with a specialized prompt that instructs it to directly generate the action, omitting the thought step. Having implicitly learned the reasoning patterns, the model can produce the correct action without the costly auto-regressive generation of the rationale text. This ”train-with-thought, infer-without-thought” paradigm allows the agent to retain the high-quality decision-making of the teacher model while operating at a fraction of the computational cost and latency.
5. Experiments
@@ 5.1.1. Datasets and Benchmarks
To evaluate real-world performance, we built MCP-Bench, a comprehensive benchmark from anonymized production logs that captures natural query distributions from simple lookups to complex multi-hop requests. Working with Baidu’s annotation team, we used a rigorous multi-stage protocol where experts annotated each query’s ground-truth MCP Servers, execution trajectories, and ideal answers. Cross-validation with multiple annotators and consensus resolution achieved a Cohen’s kappa of 0.87 for reliability.
@@ 5.1.2. Baselines
We compare TURA against a strong baseline and two ablated versions of our own system:
@@ 5.1.3. Evaluation Metrics
We employ a multi-faceted evaluation strategy.
@@
  • Online A/B Testing: In the live production environment, we track standard industry metrics: Session Success Rate(SSR)(Deng and Shi, 2016), which measures the fraction of user sessions where a satisfactory answer is provided, and Good vs. Same vs. Bad (GSB)(Zou et al., 2021), a human-rated comparison of TURA’s output against the production baseline.
@@ 5.1.4. Implementation Details
Our TURA implementation utilizes Qwen3-1.7B for query decomposition and ERNIE as the dense retrieval encoder. The DAG planner is implemented using Deepseek-V3. The agent distillation process employs Deepseek-V3 as the teacher model. The resulting student agents are fine-tuned from the Qwen3 series. For latency evaluation, 80th percentile measurements were conducted for tool execution processes. The Qwen3 series models were benchmarked on two NVIDIA L20 GPUs configured identically to the production deployment environment, while Deepseek-V3 was evaluated using the online service hosted on Baidu Qianfan platform.
@@ 5.2.1. End-to-End Offline Evaluation
We conduct comprehensive end-to-end evaluation of TURA against a strong LLM + RAG baseline on the MCP-Bench dataset. Table 1 shows that TURA achieves substantial improvements in both answer accuracy and faithfulness across human and automated evaluations.
Table 1. End-to-end performance comparison on MCP-Bench
@@
TURA demonstrates significant performance gains over the RAG baseline. In answer accuracy, TURA achieves 87.5% versus RAG’s 65.3% in human evaluation. This substantial gain highlights the limitations of passive retrieval for complex multi-faceted queries and validates our hypothesis that active tool planning is essential for robust performance.
The improvement in faithfulness is even more pronounced. TURA achieves 96.2% faithfulness compared to RAG’s 72.4% in human evaluation. This difference stems from a fundamental architectural advantage: while RAG relies on synthesis from potentially noisy text corpora and is prone to hallucination, TURA’s framework enables dynamic invocation of verified tools that provide high-fidelity information.
The strong correlation between human and LLM evaluations across both methods validates the reliability of automated evaluation approaches for this task domain.
@@ 5.2.2. Online Deployment and A/B Testing Results
Following promising offline results, TURA was deployed in a live A/B test against the incumbent LLM + RAG production system. We randomly sampled 103 user queries across multiple domains, with human evaluators assessing response quality on accuracy, content value, and overall satisfaction using a comprehensive evaluation framework.
As shown in Table 2, TURA delivered statistically significant improvements across key business metrics, demonstrating consistent advantages in both session satisfaction and response quality distribution.
@@
System Total Issues Advantage Rate SSR
LLM + RAG (Baseline) 66 - 55.1%
@@
The online results confirm TURA’s superiority. It increased the Session Success Rate by 8.9% and achieved an 8.7% overall performance advantage. In head-to-head comparisons, TURA was rated as ”Good” (strictly better than the baseline) in 13% of cases, while maintaining ”Satisfactory” performance in 86% of cases and reducing ”Bad” ratings to only 4%.
Our analysis revealed that TURA’s tool-calling capabilities were a key performance driver, enabling it to excel in scenarios requiring real-time data accuracy where the LLM + RAG baseline failed. For instance, the baseline showes significant temperature deviations in weather queries and major discrepancies in train schedules, whereas TURA provided precise, up-to-date information directly from authoritative sources. This superiority translated to a sharp reduction in critical failures from 9 in the baseline to just 4 in TURA. Overall, TURA reduced the total issue count by 16.7% (from 66 to 55), with consistent improvements across all categories: accuracy (-7.1%), content richness (-28.6%), and content value (-17.6%), indicating a substantial enhancement in response informativeness and reliability.
Given its robust performance improvements and consistent advantages across multiple evaluation metrics, TURA has demonstrated clear superiority over traditional LLM + RAG baselines, validating the effectiveness of synergizing RAG with agentic tool-use for accessing both static and dynamic information in industrial AI search productions.
@@ 5.3.1. Analysis of Intent-Aware MCP Server Retrieval (RQ2)
Figure 3. Impact of the number of synthetic queries ($N_{Q}$) per server on retrieval Recall@5. Performance peaks at 20 queries.
Next, we explored how to structure these queries in the index (Table 5). A Single-Vector approach, which concatenates all text into one document for embedding, suffers from semantic dilution and performs worst. In contrast, Multi-Vector approaches, which create separate embeddings for different parts of the server’s information, achieve superior performance. This is because they offer higher representational granularity, providing focused semantic targets. While using only synthetic queries (Queries Only) performs marginally best, we chose the Queries + Doc strategy. This method retains the original server document as a ”safety net,” ensuring robustness for queries not covered by the synthetic data, a crucial feature for real-world deployment.
@@ 5.3.2. Importance of the DAG-based Task Planner (RQ3)
While Table 1 showes the latency impact of the DAG planner on the entire dataset, we conduct a targeted analysis on a challenging subset of MCP-Bench containing only complex, multi-hop queries where parallelism is possible. This isolates the planner’s contribution to efficiency.
As Table 6 illustrates, the DAG-based planner reduces average latency by 44.2% on these complex queries by identifying and executing independent sub-tasks in parallel. This substantial efficiency gain is achieved with no degradation in the execution success rate, confirming the effectiveness of our planner in optimizing complex workflows for online latency.
@@ 5.3.3. Effectiveness of Agent Distillation (RQ4)
The empirical results, presented in Table 7, unequivocally demonstrate the efficacy of our approach. We highlight two primary findings. First, our distilled models achieve a remarkable level of performance, surpassing even the powerful teacher model. Specifically, the Qwen3-4B Distilled and Qwen3-30B-A3B Distilled models attain accuracies of 88.3% and 88.7% respectively. These results are substantially higher than both the 671B parameter teacher (82.4%) and the formidable GPT-4o baseline (81.7%). This phenomenon, where the student outperforms the teacher, validates the high quality of the synthetic trajectories generated by our data curation pipeline, which effectively filters noise and crystallizes optimal reasoning paths into a targeted training dataset.
@@ 6. Conclusion
This paper introduced TURA, a novel agentic framework designed to bridge the gap between traditional static RAG systems and the growing demand for dynamic, real-time information access in modern AI search. TURA overcomes passive retrieval’s limitations through a cohesive three-stage architecture: Intent-Aware Retrieval for precise tool selection, DAG-based Task Planning for latency-optimized parallel execution, and an efficient Distilled Agent Executor. This empowers AI search to handle complex, multi-faceted queries previously intractable for conventional RAG systems. Rigorous empirical evaluation, validated by a large-scale online A/B test in production environment, confirms TURA’s significant superiority. It markedly outperforms strong baselines, delivering substantial gains in answer accuracy and faithfulness, and a notable increase in Session Success Rate. This work presents a production-proven blueprint for the next generation of conversational AI, demonstrating a clear paradigm shift from passive information retrieval to active, tool-augmented systems. By enabling the seamless integration of heterogeneous, real-time data sources, TURA establishes a new benchmark for building robust and scalable industrial-grade AI search products.

### v2

@@ 4.2.2. Server-level Semantic Index Augmentation
A significant challenge in tool use is the “lexical gap” between user vernacular and formal API or server descriptions. To bridge this, we perform an extensive offline index augmentation process. For each server $M_{i}\in\mathcal{M}$, we first define its holistic description $D_{i}$. Then, we utilize a generative LLM, $g_{\text{LLM-gen}}$ with a specific prompt (shown in Appendix A.1), to produce a large, diverse set of synthetic queries, $\mathcal{Q}^{\text{syn}}_{i}$, that a user might plausibly issue to access the functionalities of $M_{i}$. The process can be formulated as
(3) $\mathcal{Q}^{\text{syn}}_{i}=g_{\text{LLM-gen}}(D_{i},\text{temperature}=T_{\text{high}}).$
@@ 4.3. DAG-based Task Planner
The planner, implemented with a highly capable LLM $p_{\text{LLM-plan}}$, is prompted to act as a solution architect with a specific prompt (shown in Appendix A.3). It analyzes the relationships between sub-queries and the capabilities of the retrieved servers to construct a DAG, $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
@@ 4.4.1. Trajectory Synthesis and Data Curation
This raw data is then subjected to a rigorous, automated curation pipeline. The first stage, Correctness Filtering, employs a judge model, $J_{\text{correct}}$(Huang et al., 2024), to validate each step of a trajectory. This judge scrutinizes for adherence to API schemas, validity of parameter values, and logical soundness of the thought process leading to the action. Any trajectory failing these checks is discarded.
Subsequently, the second stage, Efficiency Filtering, uses another judge, $J_{\text{efficient}}$, to analyze the now-correct trajectories for performance. It identifies and flags issues such as action redundancy and path sub-optimality These inefficient trajectories are then either pruned or programmatically corrected. This two-stage curation transforms the noisy expert data into a high-quality, optimal distillation dataset $D_{\text{distill}}$:
@@
4.4.2. Mixed-Rationale Supervised Fine-Tuning
To achieve minimal inference latency, we fine-tune Qwen3 (Yang et al., 2025) series, a much smaller model than large-scale LLMs like Deepseek-V3 on $\mathcal{D}_{\text{distill}}$ using a mixed-rationale supervised fine-tuning (SFT) strategy. The training process explicitly leverages the chain-of-thought data. The agent $\mathcal{A}_{\theta}$ is trained to predict the full sequence of tokens, including both the thought and the action. The loss function is the standard cross-entropy loss over the target sequence:
(8) $\mathcal{L}_{\text{SFT}}(\theta)=-\sum_{(st^{\prime},\pi^{\prime})\in\mathcal{D}_{\text{distill}}}\sum_{t=1}^{|\pi^{\prime}|}\log P_{\theta}(y_{t}|y_{<t},st^{\prime}),$
@@ 4.4.2. Mixed-Rationale Supervised Fine-Tuning
Critically, during online inference, we provide the agent $\mathcal{A}_{\theta}$ with a specialized prompt that instructs it to directly generate the action, omitting the thought step. Having implicitly learned the reasoning patterns, the model can produce the correct action without the costly auto-regressive generation of the rationale text. This “train-with-thought, infer-without-thought” paradigm allows the agent to retain the high-quality decision-making of the teacher model while operating at a fraction of the computational cost and latency.
5. Evaluation
@@ 5.1.1. Datasets and Benchmarks
To evaluate real-world performance, we built MCP-Bench, a comprehensive benchmark from anonymized production logs that captures natural query distributions from simple lookups to complex multi-hop requests. Working with Baidu’s annotation team, we used a rigorous multi-stage protocol where experts annotated each query’s ground-truth MCP Servers, execution trajectories, and ideal answers. Cross-validation with multiple annotators and consensus resolution achieved a Cohen’s kappa of 0.87 for reliability. MCP-Bench is constructed from anonymized production logs with a multi-stage expert annotation protocol. We report key statistics to support reproducibility:
  • Queries: 10,683 total queries, with 95.02% single-intent and 4.98% multi-intent queries.
  • MCP servers: 61 servers across 10 categories, including travel, weather, food, finance, and general-purpose “other” categories.
  • Tool-call complexity: average 1.533 tool calls per query (P50=1, P90=3), and average DAG width 1.147.
  • Domain distribution. In addition to search, which accounts for the largest proportion, the remaining top categories by tool-call proportion are travel (29.60%), finance (11.42%), weather (7.70%), food (7.22%), and other (28.57%).
  • Evaluation protocol. We release annotation prompts, judge rubrics, and aggregation scripts to facilitate reproducibility, while anonymizing server and tool identifiers to preserve privacy.
MCP-Bench is publicly released at https://github.com/weixiaolong94-hub/TURA.
@@ 5.1.2. Baselines
We compare TURA against three categories of strong baselines and two ablated versions of our own system:
@@
  • Dynamic RAG adapts retrieval decisions during generation (e.g., when/what to retrieve) but remains constrained to a static index; it cannot interact with transactional APIs.
  • Tool-Agent (ReAct-style agents and DAG-enabled agents) can call tools, but do not incorporate our unified MCP server retrieval module or our distilled low-latency executor; we run them with the same MCP servers to isolate orchestration and execution differences.
@@ 5.1.3. Evaluation Metrics
To evaluate the performance of TURA and baselines, we take a multi-faceted evaluation strategy.
@@
  • Online A/B Testing: In the live production environment, we track standard industry metrics: Session Success Rate (SSR) (Deng and Shi, 2016), which measures the fraction of user sessions where a satisfactory answer is provided, and Good vs. Same vs. Bad (GSB) (Li et al., 2025c), a human-rated comparison of TURA’s output against the production baseline.
@@ 5.1.4. Implementation Details
Our TURA implementation utilizes Qwen3-1.7B for query decomposition and ERNIE as the dense retrieval encoder. The DAG planner is implemented using Deepseek-V3. The agent distillation process employs Deepseek-V3 as the teacher model. The resulting student agents are fine-tuned from the Qwen3 series. For latency evaluation, 80th percentile measurements were conducted for tool execution processes. The Qwen3 series models were benchmarked on two NVIDIA L20 GPUs configured identically to the production deployment environment, while Deepseek-V3 was evaluated using the online service hosted on the Baidu Qianfan platform.
@@ 5.2.1. End-to-End Offline Evaluation
We conduct a comprehensive end-to-end evaluation of TURA against three strong baselines: LLM + RAG, Dynamic RAG, and Tool-Agent on the MCP-Bench dataset. Table 1 shows that TURA achieves substantial improvements in both answer accuracy and faithfulness across human and automated evaluations.
Table 1. End-to-end performance comparison on MCP-Bench.
@@
TURA demonstrates significant performance gains over three strong baselines. In answer accuracy, TURA achieves 87.5% versus RAG’s 65.3% in human evaluation. This substantial gain highlights the limitations of passive retrieval for complex multi-faceted queries and validates our hypothesis that active tool planning is essential for robust performance. The improvement in faithfulness is even more pronounced. TURA achieves 96.2% faithfulness compared to RAG’s 72.4% in human evaluation. This difference stems from a fundamental architectural advantage: while RAG relies on synthesis from potentially noisy text corpora and is prone to hallucination, TURA’s framework enables dynamic invocation of verified tools that provide high-fidelity information. The strong correlation between human and LLM evaluations across both methods validates the reliability of automated evaluation approaches for this task domain.
@@ 5.2.2. Online Deployment and A/B Testing Results
Following the promising offline result, TURA was deployed in a live A/B test against the incumbent LLM + RAG production system. We sampled 103 queries to balance coverage and cost, combining high-frequency production requests with random samples to approximate real user demand, while accounting for potential execution failures. This domain mix aligns well with MCP-Bench. We observe that our largest gains stem from travel-related scenarios, a domain characterizing real-time and constraint-heavy tasks such as itinerary planning and availability checks.
Each query is independently evaluated by trained annotators following a shared rubric. Particularly, we define accuracy as whether the response correctly satisfies the user intent, content value as the degree to which the response provides actionable and decision-supporting information, and satisfaction as the overall user-perceived utility considering correctness, completeness, and presentation. Each result is annotated by three annotators, and in cases of disagreement, the final decision is made through voting.
As shown in Table 2, TURA delivered statistically significant improvements across key business metrics, demonstrating consistent advantages in both session satisfaction and response quality distribution. The online results confirm TURA’s superiority. It increased the Session Success Rate by 8.9% and achieved an 8.7% overall performance advantage. In head-to-head comparisons, TURA was rated as ”Good” (strictly better than the baseline) in 13% of cases, while maintaining ”Satisfactory” performance in 86% of cases and reducing ”Bad” ratings to only 4%.
@@
Online System Total Issues Advantage Rate SSR
LLM + RAG (Legacy) 66 - 55.1%
@@
Table 3 presents the detailed issue statistics and comparative performance metrics between TURA and the LLM + RAG (i.e., online legacy system). The comparative results revealed that TURA’s tool-calling capabilities were a key performance driver, enabling it to excel in scenarios requiring real-time data accuracy where the LLM + RAG baseline failed. For instance, the baseline shows significant temperature deviations in weather queries and major discrepancies in train schedules, whereas TURA provided precise, up-to-date information directly from authoritative sources. This superiority translated to a sharp reduction in critical failures from 9 in the baseline to just 4 in TURA. Overall, TURA reduced the total issue count by 16.7% (from 66 to 55), with consistent improvements across all categories: accuracy (-7.1%), content richness (-28.6%), and content value (-17.6%), indicating a substantial enhancement in response informativeness and reliability. Given its robust performance improvements and consistent advantages across multiple evaluation metrics, TURA has demonstrated clear superiority over traditional LLM + RAG baselines, validating the effectiveness of synergizing RAG with agentic tool-use for accessing both static and dynamic information in industrial AI search productions.
@@ 5.3.1. Analysis of Intent-Aware MCP Server Retrieval (RQ2)
Figure 3. Impact of the number of synthetic queries per server on retrieval Recall@5. Performance peaks at 20 queries.
@@
Next, we explored how to structure these queries in the index (Table 5). A Single-Vector approach, which concatenates all text into one document for embedding, suffers from semantic dilution and performs the worst. In contrast, Multi-Vector approaches, which create separate embeddings for different parts of the server’s information, achieve superior performance. This is because they offer higher representational granularity, providing focused semantic targets. While using only synthetic queries (Queries Only) performs marginally best, we chose the Queries + Doc strategy. This method retains the original server document as a “safety net”, ensuring robustness for queries not covered by the synthetic data, a crucial feature for real-world deployment.
@@ 5.3.2. Importance of the DAG-based Task Planner (RQ3)
While Table 1 shows the latency impact of the DAG planner on the entire dataset, we conduct a targeted analysis on a challenging subset of MCP-Bench containing only complex, multi-hop queries where parallelism is possible. This isolates the planner’s contribution to efficiency. As Table 6 illustrates, the DAG-based planner reduces average latency by 44.2% on these complex queries by identifying and executing independent sub-tasks in parallel. This substantial efficiency gain is achieved with no degradation in the execution success rate, confirming the effectiveness of our planner in optimizing complex workflows for online latency.
@@ 5.3.3. Effectiveness of Agent Distillation (RQ4)
The empirical results, presented in Table 7, unequivocally demonstrate the efficacy of our approach. We highlight two primary findings. First, our distilled models achieve a remarkable level of performance, surpassing even the powerful teacher model. Specifically, the Qwen3-4B Distilled and Qwen3-30B-A3B Distilled models attain accuracies of 88.3% and 88.7%, respectively. These results are substantially higher than both the 671B parameter teacher (82.4%) and the formidable GPT-4o baseline (81.7%). This phenomenon, where the student outperforms the teacher, validates the high quality of the synthetic trajectories generated by our data curation pipeline, which effectively filters noise and crystallizes optimal reasoning paths into a targeted training dataset.
@@ 6. Conclusion
This paper introduced TURA, a novel agentic framework designed to bridge the gap between traditional static RAG systems and the growing demand for dynamic, real-time information access in modern AI search. TURA overcomes passive retrieval’s limitations through a cohesive three-stage architecture: Intent-Aware Retrieval for precise tool selection, DAG-based Task Planning for latency-optimized parallel execution, and an efficient Distilled Agent Executor. This empowers AI search to handle complex, multi-faceted queries previously intractable for conventional RAG systems. Rigorous empirical evaluation, validated by a large-scale online A/B test in a production environment, confirms TURA’s significant superiority. It markedly outperforms strong baselines, delivering substantial gains in answer accuracy and faithfulness, and a notable increase in Session Success Rate. This work presents a production-proven blueprint for the next generation of conversational AI, demonstrating a clear paradigm shift from passive information retrieval to active, tool-augmented systems. By enabling the seamless integration of heterogeneous, real-time data sources, TURA establishes a new benchmark for building robust and scalable industrial-grade AI search products.